OpenStudy (anonymous):

-5over3x = -4over3? show work

OpenStudy (saifoo.khan):

\[\frac{-5}{3x} = \frac{-4}{3}\]

OpenStudy (saifoo.khan):

cancel both -ve signs.

OpenStudy (anonymous):

then multiply by 3

OpenStudy (anonymous):

and x too

OpenStudy (anonymous):

then divide by 4

OpenStudy (anonymous):

\[15=12x\] people love to "cross multiply"!

OpenStudy (saifoo.khan):

\[\frac{5}{3x}= \frac{4}{3}\] Cross multiply. \[3x*4 = 5*3 \]12x = 15 \[x = 15/12 \]x = 5/4

OpenStudy (anonymous):

\[\frac{15}{12}=\frac{5}{4}=x\]

OpenStudy (anonymous):

saifooooo! you awake tonight?